OpenStudy (anonymous):

I'm not sure how to do this :((((( The temperature outside is 2° Celsius. The formula for degrees Celsius is: C = (F - 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 2 degrees Celsius is.

OpenStudy (anonymous):

actually the conversion equation is this: \[C = {5 \over 9}(F-32)\] If you're book really said what you posted use that. But either way lets start off by trying to solve for F. do you know how to do that?

OpenStudy (anonymous):

Oh yeah thanks and no :( how?

OpenStudy (anonymous):

Lets first multiply 9 to the other side \[9*C=5(F-32)\] Now divide by 5 \[{9 \over 5}C=F -32\] Can you solve the rest?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

how do you think we would get 32 on the other side of the equation?

OpenStudy (anonymous):

uhmmmm +32?

OpenStudy (anonymous):

Good job! that's correct. So did my first few steps to get there make sense?

OpenStudy (anonymous):

well I didn't get hoy you got the 9 over 5 :(

OpenStudy (anonymous):

I don't know what to do now :(((

OpenStudy (anonymous):

multiply Celsius with 9/5 and then add 32 you will have Fahrenheit and multiply F by 5/9 and then subtract 32 you will have Celsius

OpenStudy (anonymous):

oae pammi u got it or not?? @pam142